# Improved IR-based Bug Localization with Intelligent Relevance Feedback

### v1

@@ Abstract
Bug Localization, Query Reformulation, Intelligent Relevance Feedback, Information Retrieval, Large Language Models, Natural Language Processing, Software Engineering
@@ I Introduction
  • A replication package with a prototype, a curated dataset, and configuration details for third-party replication and reuse.
*https://shorturl.at/Y8xqi
@@ III-B Intelligent Relevance Feedback
Once we have the results from Elasticsearch, we employ advanced prompt methods and Large Language Models (LLM) to determine the relevance between a bug report and each result (i.e., source code). LLMs have shown remarkable capabilities understanding natural language texts and source code[41, 42, 43]. We leverage their capabilities to capture intelligent relevance feedback against a query (a.k.a., bug report) as follows. To achieve this, we use prompt engineering, document segmentation, and finally relevance estimation as follows.
Prompt Engineering: Prompting is a novel method that instructs the LLMs (e.g., Mistral) to generate meaningful responses without any expensive training [44, 45, 46]. It involves crafting appropriate instructions to guide LLM outputs and make them applicable to different problem-solving tasks[47, 48, 49, 50, 51]. LLMs have been found to be effective with well-designed prompts that are clear, specific, and actionable [45, 52]. We first developed a candidate prompt to determine whether a given code segment triggers a reported bug. It instructs the LLM to find the relevance, deliver the output in a JSON format, and act as a rational software engineer, incorporating the contextual information from the bug report and code segment.
To refine our candidate prompt, we employed SAMMO [53], a framework that explores and evaluates various prompt configurations using beam search. We configured SAMMO with LLaMA-3[54] and used a small dataset of 20 bug reports with corresponding buggy code segments (ground truth) to guide our optimization process. SAMMO iteratively generated prompt variants by applying various modification operations to the candidate prompt with LLaMA’s assistance. In each iteration, we used LLaMA, bug reports and ground truth code to determine the fitness of each prompt and provide a performance update to SAMMO. Through an extensive search, SAMMO was able to find an optimized version of the prompt. Table II shows the optimized prompt template, used in the subsequent steps of our technique. Both candidate and optimized prompts can be found in the replication package.
@@ III-C Query Expansion
Here, $PR(V_{i})$ represents the PageRank of vertex $V_{i}$. The term $d$ denotes the damping factor with a default set to 0.85. $N$ refers to the total number of vertices in the graph, while $M(V_{i})$ indicates the set of vertices linked to $V_{i}$. Finally, $PR(V_{j})$ and $L(V_{j})$ represent the PageRank and outbound links of vertex $V_{j}$.
@@ III-D2 Rescoring
Finally, we rank the source documents based on their scores for their potential to be buggy (Step 9, Fig. 2) and return the top-K (e.g., K=10) documents. Our scoring process aims to bridge the gap between bug reports and source code by incorporating a deeper contextual understanding of the LLM and going beyond their textual and semantic relevance (Table I).
@@ IV-A Dataset Construction
In our experiment, we used the Bench4BL [21], a comprehensive benchmark dataset that contains 10,017 bug reports from 51 open-source systems, covering a total of 695 software versions. Our initial assessment revealed that bug reports from the older systems lacked crucial versioning information, making them unsuitable for our study. Additionally, we could not accurately link some bug reports to their corresponding buggy code within the code repositories. Hence, we excluded these bug reports from our dataset. We also found bug reports containing only stack traces without accompanying any textual descriptions of their bugs. We identified those bug reports using regular expressions[10] and excluded them from the dataset. Following these refinement steps, our final dataset comprised 4,683 bug reports from 42 different systems spanning across 684 versions. Table III-a summarizes our curated dataset.
@@ IV-B Evaluation Metrics
Here, $MRR$ averages the Reciprocal Ranks ($RR_{q}$) across all queries $q$ in set $Q$, where $RR_{q}$ is the Reciprocal Rank for query $q$.
@@
Changed Documents Bug Report Count Baseline LLaMA Mistral Qwen
@@ IV-D Evaluating BRaIn
According to our investigation, some bugs trigger changes to a single document during bug resolution, whereas others trigger changes to multiple documents. We thus evaluate BRaIn’s performance in localizing bugs that warrant changes across multiple source documents. Table V shows the performance of BRaIn, powered by three different LLMs, in terms of MAP, MRR, and HIT@10. We grouped bugs from our dataset into four categories based on the number of their changed documents: 1, 2, 3, and 4 or more. Our findings show that BRaIn performs strongest when paired with Mistral. For the 1,949 bugs requiring changes to a single document, BRaIn (Mistral) achieves a MAP of 0.542, representing a significant 14.3% improvement over the baseline counterpart. The improvements extend to other metrics, with a 14.3% increase in MRR and 7.1% in HIT@10. BRaIn (Mistral) also excels in resolving bugs that require multiple document changes, achieving improvements of 7.0-10.6% in MAP, 7.9-10.6% in MRR, and 6.0-6.9% in HIT@10. The other variants of BRaIn also outperform the baseline in localizing bugs that require multiple document changes, achieving improvements of up to 9.2% in MAP, 10.0% in MRR, and 5.1% in HIT@10.
@@
Baseline BLUiR Blizzard Rocchio Sysman-SCP BRaIn
@@
To determine the impact of query expansion in isolation, we evaluated BRaIn’s performance without the reranking component (Table VI). Interestingly, all BRaIn variants achieved similar MAP scores of around 0.49, falling short of optimal performance. For example, BRaIn (LLaMA) and BRaIn (Mistral) saw MAP scores decrease by 7.2% and 7.8%, respectively, while their MRR scores remained relatively stable. BRaIn (Qwen), on the other hand, showed a 7.8% increase in MRR alongside a slight improvement in MAP. However, all BRaIn variants outperformed the baseline by 10.5-10.9% in MRR and 2.4-2.9% in MAP.
In contrast, when reranking was applied isolately, MAP scores for BRaIn (LLaMA), BRaIn (Mistral), and BRaIn (Qwen) dropped by 3.0%, 3.2%, and 6.1%, respectively, while their MRR scores remained consistent. Despite these decreases, all variants showed improvements of 3.1-11% in MRR and 1.0-7.1% in MAP over the baseline.
These findings show that Intelligent Relevance Feedback improves the individual components of query expansion and reranking. However, they also underscore the importance of the synergy between these components in achieving optimal performance.
@@
To replicate the traditional IR-based Baseline VSM, we index all source documents of a repository and use bug reports (title + description) as queries. These queries are executed with the Elasticsearch [40], which retrieves relevant documents using the BM25 algorithm [39] and Boolean search, with default parameters for k and b. Other traditional approaches from literature– BLUiR [35] and Blizzard [10]– use structured information from bug reports and source code for bug localization. BLUiR calculates suspiciousness scores using class names, method names, variable names, comments, and bug report elements (title, description), combining multiple searches into an overall score. Blizzard categorizes bug reports into three types and constructs text graphs from these reports to generate queries and retrieve relevant buggy source documents. For both approaches, we employ Apache Lucene[68] for retrieval. We replicated these methods by adapting them from Bench4BL repository[21] and Blizzard’s replication package[69] from the authors. We compare BRaIn with these IR-based techniques to assess the performance of our technique against established models in the field.
@@
Since Machine Learning (ML) techniques can capture complex patterns in data using non-linear relationships, we compare BRaIn against three ML-based techniques– DNNLOC [36], NextBug [37], and RLocator [38]. DNNLOC combines multiple features— rVSM score[71] for bug report-source code similarity, class name similarity, collaborative filtering, and bug report recency and frequency—and uses a neural network to predict suspiciousness scores to rank documents. NextBug employs Word2Vec [72] embeddings to capture semantic relations between bug reports and source code and thus to localize the buggy documents. In our experiments, we substituted Word2Vec with CodeT5 embeddings [73] to capture more nuanced text-level semantic associations, as opposed to token-level. RLocator is a recent deep-learning technique that employs a reinforcement learning model, framed as a Markov Decision Process, to optimize ranking of buggy documents. We replicated DNNLOC and NextBug by following the respective authors’ approaches and replicate RLocator using the authors’ provided replication package on Zenodo[74].
Table VII-a summarizes our comparison details with the baseline techniques. Among traditional IR-based approaches, Blizzard achieves a MAP score of 0.506, while Baseline VSM and BLUiR scores are 0.484 and 0.450. BRaIn outperforms them with a MAP of 0.537, achieving a maximum improvement of 19.3% over these techniques. Similarly, BRaIn achieves notable gains in MRR and HIT@10, with increases of up to 17.5% and 12.2%. Among the IR based approaches that leverage relevance feedback, Rocchio’s algorithm achieves a MAP of 0.489, slightly above the baseline, while Sysman-SCP falls short by 2.5%. BRaIn again leads here with the improvements in MAP, MRR, and HIT@10 of 13.8%, 5.5%, and 3.7%, respectively. These results underscore the advantages of Intelligent Relevance Feedback, which uses contextual understanding over traditional techniques based on textual relevance for bug localization.
@@ V-A IR based Bug Localization
In our work, we use a VSM-based retrieval that is enhanced by Boolean search (e.g., Elasticsearch). However, it was complemented by contextual understanding of bugs and Intelligent Relevance Feedback, leveraging the capabilities of LLMs.
@@ V-B Query Reformulation
Poorly constructed queries from software bug reports can significantly hinder IR-based bug localization[82, 11]. To tackle this issue, researchers have developed query reformulation techniques that can improve search queries by incorporating better terms or eliminating unnecessary ones. For instance, Refoqus[29] uses query characteristics and machine learning to recommend strategies like query reduction or expansion for a given query. Graph-based methods analyze semantic and syntactic relationships within bug reports to identify key terms. Rahman and Roy[10] create text graphs to collect important terms based on three different bug types to improve query reformulation. The authors later demonstrate the presence of optimal queries in bug reports by employing Genetic algorithms[11]. Relevance feedback-based approaches iteratively refine queries based on their search results. Gay et al.[29] used Rocchio’s algorithm to improve queries with developer feedback, while Sisman et al.[30] expanded queries by selecting terms from the top-ranked documents using Spatial Code Proximity (SCP), without using any explicit relevance feedback.
@@ V-C Deep Learning and Bug Localization
Recent advancements in deep learning have encouraged its applications in bug localization. DNNLOC[36], a seminal work on this topic, identifies buggy documents by learning from multiple text-based features and metadata (e.g., rVSM score[71], class name similarity, bug report recency). However, its reliance on features like bug fixing recency can limit its application [71]. A recent technique, FBL-BERT [83] uses a BERT-based model, ColBERT [84], for document scoring with late interaction. However, it relies on changesets for resolution, which are difficult to track in large, fast-changing projects. Another recent technique, RLocator[38], aims to optimize ranking metrics in the bug localization process. It formulates bug localization as a Markov Decision Process (MDP) and employs reinforcement learning (RL) to guide its decisions. Other approaches like TRANP-CNN [85] and CooBa [86] use Convolutional Neural Networks (CNN) and Graph Convolutional Networks (GCN) to improve cross-project bug localization. However, these methods may suffer from a lack of scalability when handling large volumes of documents.
In contrast, our technique deals with a limited set of documents for scalability, retrieved by a scalable tool – Elasticsearch[40]. By combining efficient IR-based filtering with the contextual depth of language models, BRaIn accurately narrows down and ranks buggy documents, improving bug localization.
@@ VI Threats to Validity
Threats to construct validity concern the appropriateness of our evaluation metrics. We employed widely used metrics such as Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K, which were commonly used in existing literature on bug localization [10, 35, 87] and Information Retrieval studies[88, 89]. Therefore, this choice of metrics minimizes threats to construct validity.
Finally, we used 20 bug reports from our dataset to optimize prompts with LLaMA. Since they are a part of our experimental dataset, it could introduce bias. We repeated a limited experiment and found similar performance to the reported ones in the paper. Thus, any relevant threat of bias is minimal to none.
@@ VII Conclusion and Future Work
Software bugs consume resources and cause financial losses[4]. Resolving these challenges has been a major focus, with researchers working on solutions for decades. In this paper, we introduced BRaIn to address the contextual gaps between bug reports and source codes by assessing the relevance between bug reports and code with Large Language Models (LLM). Our proposed technique leverages the LLM’s feedback (a.k.a., Intelligent Relevance Feedback) to improve bug localization by reformulating queries and reranking source documents. We evaluated BRaIn’s performance using three widely used metrics in bug localization: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. BRaIn achieved substantial improvements, showing increases of 87.6% in MAP, 89.5% in MRR, and 48.8% in HIT@K compared to baseline techniques. Furthermore, BRaIn detected $\approx$52more low-quality bugs within the top 10 ranks relative to the baseline VSM.

### v2

@@ Abstract
Bug Localization, Intelligent Relevance Feedback, Query Reformulation, Information Retrieval, LLM
@@ I Introduction
  • A replication package[39] with a prototype, a curated dataset, and configuration details for third-party replication and reuse.
@@ III-B Intelligent Relevance Feedback
Once we have the results from Elasticsearch, we employ advanced prompt methods and Large Language Models (LLM) to determine the relevance between a bug report and each result (i.e., source code). LLMs have shown remarkable capabilities understanding natural language texts and source code[42, 43, 44]. We leverage their capabilities to capture intelligent relevance feedback against a query (a.k.a., bug report). To achieve this, we use prompt engineering, document segmentation, and finally relevance estimation as follows.
Prompt Engineering: Prompting is a novel method that instructs the LLMs (e.g., Mistral) to generate meaningful responses without any expensive training [45, 46, 47]. It involves crafting appropriate instructions to guide LLM outputs and make them applicable to different problem-solving tasks[48, 49, 50, 51, 52]. LLMs have been found to be effective with well-designed prompts that are clear, specific, and actionable [46, 53]. Following the insights, we first developed a candidate prompt based on efficient prompt-building guidelines[54, 55]. Our goal was to determine whether a given code segment triggers a reported bug. It instructs the LLM to find the relevance, deliver the output in a JSON format, and act as a rational software engineer, incorporating the contextual information from the bug report and code segment.
To refine our candidate prompt, we employed SAMMO [56], a compile-time framework that optimizes prompts by exploring various configurations through mutation operations. We configured SAMMO with LLaMA-3[57] and used a small dataset of 20 bug reports with corresponding buggy code segments (ground truth) to guide our optimization process. SAMMO iteratively generated prompt variants by applying various modification operations to the candidate prompt with LLaMA’s assistance. In each iteration, we used LLaMA, bug reports and ground truth code to determine the fitness of each prompt and provide a performance update to SAMMO. Through an extensive search, SAMMO was able to find an optimized version of the prompt. Table II shows the optimized prompt template, used in the subsequent steps of our technique.
@@ III-C Query Expansion
Here, $PR(V_{i})$ is the PageRank of vertex $V_{i}$. The term $d$ denotes the damping factor with a default set to 0.85. $N$ is the total number of vertices, while $M(V_{i})$ is the set of vertices linked to $V_{i}$. Finally, $PR(V_{j})$ and $L(V_{j})$ denote the PageRank and outbound links of vertex $V_{j}$.
@@ III-D2 Rescoring
Finally, we rank the source documents based on their scores for their potential to be buggy (Step 9, Fig. 2) and return the top-K (e.g., K=10) documents. Our scoring process aims to bridge the gap between bug reports and source code by incorporating deeper contextual understanding of the LLM beyond textual and semantic relevance (Table I).
@@ IV-A Dataset Construction
In our experiment, we used the Bench4BL [21], a comprehensive benchmark dataset that contains 10,017 bug reports from 51 open-source systems, covering a total of 695 software versions. Our initial assessment revealed that bug reports from the older systems lacked crucial versioning information, making them unsuitable for our study. Additionally, we could not accurately link some bug reports to their corresponding buggy code within the code repositories. Hence, we excluded these bug reports from our dataset. We also found bug reports containing only stack traces without any textual descriptions of their bugs. We identified those bug reports using regular expressions[10] and excluded them from the dataset. Following these refinement steps, our final dataset comprised 4,683 bug reports from 42 different systems spanning across 684 versions. Table III-a summarizes our curated dataset.
@@ IV-B Evaluation Metrics
Here, $MRR$ averages the Reciprocal Ranks ($RR_{q}$) across all queries $q$ in set $Q$.
@@ IV-C Selection of LLM
Changed Documents Bug Report Count Elasticsearch (ES) LLaMA Mistral Qwen
@@ IV-D Evaluating BRaIn
According to our investigation, some bugs trigger changes to a single document during bug resolution, whereas others trigger changes to multiple documents. We thus evaluate BRaIn’s performance in localizing bugs that warrant changes across multiple source documents. Table V shows the performance of BRaIn, powered by three different LLMs, in terms of MAP, MRR, and HIT@10. We grouped bugs from our dataset into four categories based on the number of their changed documents: 1, 2, 3, and 4 or more. Our findings show that BRaIn performs strongest when paired with Mistral. For the 1,949 bugs requiring changes to a single document, BRaIn (Mistral) achieves a MAP of 0.542, representing a significant 14.3% improvement over the baseline counterpart. The improvements extend to other metrics, with a 14.3% increase in MRR and 7.1% in HIT@10. BRaIn (Mistral) also excels in resolving bugs that require multiple document changes, achieving improvements of 7.0-10.6% in MAP, 7.9-10.6% in MRR, and 6.0-6.9% in HIT@10. The other variants of BRaIn also outperform the Elasticsearch baseline in localizing bugs that require multiple document changes, achieving improvements of up to 9.2% in MAP, 10.0% in MRR, and 5.1% in HIT@10.
@@
ES BLUiR Blizzard Rocchio Sysman-SCP BRaIn
@@
To determine the impact of query expansion in isolation, we evaluated BRaIn’s performance without the reranking component (Table VI). Interestingly, all BRaIn variants achieved similar MAP scores of around 0.49, falling short of optimal performance. For example, BRaIn (LLaMA) and BRaIn (Mistral) saw MAP scores decrease by 7.2% and 7.8%, respectively, while their MRR scores remained relatively stable. BRaIn (Qwen), on the other hand, showed a 7.8% increase in MRR alongside a slight improvement in MAP. However, all BRaIn variants outperformed the Elasticsearch baseline by 10.5-10.9% in MRR and 2.4-2.9% in MAP.
In contrast, when reranking was applied isolately, MAP scores for BRaIn (LLaMA), BRaIn (Mistral), and BRaIn (Qwen) dropped by 3.0%, 3.2%, and 6.1%, respectively, while their MRR scores remained consistent. Despite these decreases, all variants showed improvements of 3.1-11% in MRR and 1.0-7.1% in MAP over the Elasticsearch baseline.
These findings show how Intelligent Relevance Feedback improves query expansion and reranking. However, they also underscore the importance of the synergy between these components for optimal performance.
@@
To replicate the traditional IR-based bug localization with Elasticsearch baseline, we index all source documents of a repository and use bug reports (title + description) as queries. These queries are executed with Elasticsearch [41], which retrieves relevant documents using the BM25 algorithm [40] and Boolean queries, with default parameters for k and b. Other traditional approaches from literature– BLUiR [35] and Blizzard [10]– use structured information from bug reports and source code for bug localization. BLUiR calculates suspiciousness scores using class names, method names, variable names, comments, and bug report elements (title, description), combining multiple searches into an overall score. Blizzard categorizes bug reports into three types and constructs text graphs from these reports to generate queries and retrieve relevant buggy source documents. For both approaches, we employ Apache Lucene[71] for retrieval. We replicated these methods by adapting them from Bench4BL repository[21] and Blizzard’s replication package[72] from the authors. We compare BRaIn with these established IR-based techniques to validate our technique and place it in the literature.
@@
Since Machine Learning (ML) techniques can capture complex patterns in data using non-linear relationships, we compare BRaIn against three ML-based techniques– DNNLOC [36], NextBug [37], and RLocator [38]. DNNLOC combines multiple features— rVSM score[74] for bug report-source code similarity, class name similarity, collaborative filtering, and bug report recency and frequency—and uses a neural network to predict suspiciousness scores to rank documents. NextBug employs Word2Vec [75] embeddings to capture semantic relations between bug reports and source code and thus to localize the buggy documents. In our experiments, we substituted Word2Vec with CodeT5 embeddings [76] to capture more nuanced text-level semantic associations, as opposed to token-level. RLocator is a recent deep-learning technique that employs a reinforcement learning model, framed as a Markov Decision Process, to optimize ranking of buggy documents. We replicated DNNLOC and NextBug by following the respective authors’ approaches and replicated RLocator using the authors’ provided replication package on Zenodo[77]. To ensure consistency with the authors’ specifications, we replicated the methods using cross-validation.
Table VII-a summarizes our comparison details with the baseline techniques. Among traditional IR-based approaches, Blizzard achieves a MAP score of 0.506, while Elasticsearch (ES) and BLUiR scores are 0.484 and 0.450. BRaIn outperforms them with a MAP of 0.537, achieving a maximum improvement of 19.3% over these techniques. Similarly, BRaIn achieves notable gains in MRR and HIT@10, with increases of up to 17.5% and 12.2%. Among the IR based approaches that leverage relevance feedback, Rocchio’s algorithm achieves a MAP of 0.489, slightly above the baseline, while Sysman-SCP falls short by 2.5%. BRaIn again leads here with the improvements in MAP, MRR, and HIT@10 of 13.8%, 5.5%, and 3.7%, respectively. These results underscore the advantages of Intelligent Relevance Feedback, which uses contextual understanding over traditional techniques based on textual relevance for bug localization.
@@ V-A IR based Bug Localization
In our work, we use a textual similarity-based retrieval with BM25 in Elasticsearch. However, it was complemented by contextual understanding of bugs and Intelligent Relevance Feedback, leveraging the capabilities of LLMs.
@@ V-B Query Reformulation
Poorly constructed queries from software bug reports can significantly hinder IR-based bug localization[85, 11]. To tackle this issue, researchers have developed query reformulation techniques that can improve search queries by incorporating better terms or eliminating unnecessary ones. For instance, Refoqus[29] uses query characteristics and machine learning to recommend strategies like query reduction or expansion for a given query. Graph-based methods analyze semantic and syntactic relationships within bug reports to identify key terms. Rahman and Roy[10] create text graphs to collect important terms based on three different bug types to improve query reformulation. The authors later demonstrate generating optimal queries in bug reports by employing Genetic algorithms[11], which iteratively refine queries based on search results. Gay et al.[29] used Rocchio’s algorithm to improve queries with developer feedback, while Sisman et al.[30] expanded queries by selecting terms from the top-ranked documents using Spatial Code Proximity (SCP), without using any explicit relevance feedback.
@@ V-C Deep Learning and Bug Localization
Recent advancements in deep learning have encouraged its applications in bug localization. DNNLOC[36], a seminal work on this topic, identifies buggy documents by learning from multiple text-based features and metadata (e.g., rVSM score[74], class name similarity, bug report recency). However, its reliance on features like bug fixing recency can limit its application [74]. A recent technique, FBL-BERT [86] uses a BERT-based model, ColBERT [87], for document scoring with late interaction. However, it relies on changesets for resolution, which are difficult to track in large, fast-changing projects. Another recent technique, RLocator[38], optimizes ranking metrics in the bug localization process. It formulates bug localization as a Markov Decision Process (MDP) and employs reinforcement learning (RL) to localize bugs. Other approaches like TRANP-CNN [88] and CooBa [89] use Convolutional Neural Networks (CNN) and Graph Convolutional Networks (GCN) to improve cross-project bug localization. However, these methods may struggle with scalability when handling large volumes of documents.
In contrast, our technique ensures scalability by using a limited set of documents retrieved via Elasticsearch[41]. By combining efficient IR-based filtering with the contextual depth of language models, BRaIn accurately ranks buggy documents, improving bug localization.
@@ VI Threats to Validity
Threats to construct validity concern the appropriateness of our evaluation metrics. We employed widely used metrics such as Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K, which were commonly used in existing literature on bug localization [10, 35, 90] and Information Retrieval studies[91]. Therefore, this choice of metrics minimizes threats to construct validity.
Finally, we used 20 bug reports from our dataset to optimize prompts with LLaMA. Since they are part of our experimental dataset, it could introduce bias. We repeated a limited experiment and found similar performance to the reported ones in the paper. Thus, any relevant threat of bias is minimal.
@@ VII Conclusion and Future Work
Software bugs consume resources and cause financial losses[4]. Resolving these challenges has been a major focus, with researchers working on solutions for decades. In this paper, we introduced BRaIn to address the contextual gaps between bug reports and source codes by assessing the relevance between bug reports and code with Large Language Models (LLM). Our proposed technique leverages the LLM’s feedback (a.k.a., Intelligent Relevance Feedback) to improve bug localization by reformulating queries and reranking source documents. We evaluated BRaIn’s performance using three widely used metrics in bug localization: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. BRaIn achieved substantial improvements, showing increases of 87.6% in MAP, 89.5% in MRR, and 48.8% in HIT@K compared to baseline techniques.